A woman buys 3 pairs of jeans for $99.84, including tax. If each pair of jeans cost $32, what was the sales tax?: A woman buys 3 pairs of jeans for $99.84, including tax. If each pair of jeans cost $32, what was the sales tax?
Answer by jojo14344(814)

J[1]+J[2]+J[3]=subtotal+tax=total
32+32+32=96+tax=99.84
tax=99.84-96
tax=3.84
Thank you,
Jojo